The acceleration of a point is given by a=5-t^2. write an equation for velocity if v=4 m/s when t=3 s
1
Expert's answer
2019-08-12T10:07:34-0400

Velocity of the point

"v(t)=\int a(t)dt=\int (5-t^2)dt=5t-\frac{t^3}{3}+v_0"

"v(3)=5\times 3-\frac{3^3}{3}+v_0=4"

"v_0=-2 \:\rm{m/s}"

Finally, the equation for velocity

"v(t)=5t-\frac{t^3}{3}-2"